OpenStudy (anonymous):

You are employed at a department store that offers an employee discount of 14% that may be applied to purchases. You purchase some clothes originally marked at $145 that are on sale at 40% OFF. The sales tax is 6%. What is the total amount you must pay for this purchase? (Round your answer to two decimal places.) I just am having an issue with how to work this problem

OpenStudy (anonymous):

marked price is 145 total discount=40+14=54 54%of 145=54*145/100=78.3 sales tax to be paid=6*145/100=8.7 amount to be paid=145+8.7-78.3